Please help me with this problem!
The vector v is represented by the directed line segment PQ. Write v in the form ai + bj and find ||v||. P = (-3,1) Q = (4,-2)
Thank you, Tami

Answer by Alan3354(21584)

The vector v is represented by the directed line segment PQ. Write v in the form ai + bj and find ||v||. P = (-3,1) Q = (4,-2)
------------------
Difference in x = 7, in y = -3
v = 7i - 3j
||v|| = sqrt(7^2 + 3^2)
= sqrt(58)
I'm pretty sure.